1. You are studying 3 genes X, Y, and Z that are linked (in that order) in the Imperial Scorpion Pandinus imperator. The distance between X and Y is 10 cM, the distance between Y and Z is 8 cMShow work
Answer #1

Answer:

b). It has interference.

Interference refers to the reduction in the probability of occurrence of crossover at one location on a chromosome due to crossover occurring in another location. It means that the detection of the double recombinant classes shows that double crossovers must occur.

X---------10cM---------Y---------8cM----------Z

The distance between the genes = Recombination frequency (%)

The expected double crossover frequency = (Frequency between X & Y) * (Frequency between Y & Z

= 10/100 * 8/100 = 0.1 * 0.08 = 0.008 = 0.8%

0.8% out of 1500 = 12

Observed double crossovers = 5

Coefficient of Coincidence (COC) = Observed double crossovers / Expected double crossovers

= 5/12 = 0.42

Interference = 1-COC = 1-0.42 = 0.58